Circular No. 8413

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2004et IN NGC 6946 T. Zwitter, University of Ljubljana; and U. Munari, Astronomical Ob-

servatory, Padova, report the discovery by S. Moretti, on unfiltered CCD images taken on Sept. 27 with a 0.4-m telescope at Ravenna, of a bright supernova (mag 12.8) in NGC 6946. The new object is located at α = 20^{h}35^{m}25^{s}.33, δ = +60^{o}07′17′′.7 (equinox 2000.0). A high-resolution spectrogram, secured on Sept. 28 with the échelle spectrograph at the 1.82-m telescope in Asiago, shows a relatively featureless spectrum with a very broad, low-contrast Hα emission. Sharp interstellar Na I absorption lines are observed at heliocentric radial velocities of −21.1 (± 0.7) and +45.2 (± 0.4) km/s, matching expectation for separate origin within our galaxy and NGC 6946. The equivalent width of Na I D_{2} lines corresponds to a total reddening of E(B − V) = 0.41 mag, following the calibration by Munari and Zwitter (1997, A.Ap. 318, 269).

H. Yamaoka, Kyushu University, reports that K. Itagaki (Teppo-cho, Yamagata, Japan) found nothing at the location of SN 2004et on an unfiltered CCD image taken on Sept. 19.655 UT with a 0.60-m reflector (limiting mag 18.5). A. Klotz, C. Pollas and M. Boer, Observatoire de Haute Provence, report R magnitudes for 2004et, obtained with the robotic TAROT telescope: Sept. 19.989, [19.2 ± 0.45; 21.021, [17.6 ± 1.3; 22.017, [19.4 ± 1.2; 22.983, 15.17 ± 0.16; 25.978, 12.7 ± 0.3; 27.035, 12.8 ± 0.4; 27.974, 12.2 ± 0.1; 28.092, 12.6 ± 0.5.

W. Li and A. V. Filippenko, University of California, Berkeley; and S. D. Van Dyk, Infrared Processing and Analysis Center, California Institute of Technology, report that SN 2004et was imaged on Sept. 30 UT with the 0.76-m Katzman Automatic Imaging Telescope (KAIT) at Lick Observatory, yielding position end figures 25^{s}.37, 17′′.8 (equinox 2000.0; uncertainty ∼ 0′′.2 in each coordinate). A faint object, possibly extended, appears at this location on multi-band images of NGC 6946 taken with the Kitt Peak 0.9-m telescope on 1989 May 30. Based on calibrations of the field on three photometric nights with the Lick Nickel 1-m telescope, they measure the object at R = 20.8 ± 0.2, adding that it was not detected in the Kitt Peak B- and I-band images (B > 21.3, I > 19.7). Using a distance of 5.5 Mpc to NGC 6946 (Shapley et al. 2001, Ap.J. Suppl. 137, 139), and E(B − V) = 0.41 mag (from Zwitter and Munari, above), the absolute magnitudes of the object are M_{B} > −9.1, M_{R} = −8.9, and M_{I} > −9.7. The high luminosity makes it likely to be a star cluster rather than a single star. Unfortunately, there do not appear to be Hubble Space Telescope or KAIT pre-explosion images of the site of SN 2004et.

2004 September 30  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green